# Undesirable Memorization in Large Language Models: A Survey

### v1

@@ III Measuring memorization
Measuring memorization in large language models is a complex task, primarily because we cannot easily compute the probability of all the samples present in the training data. To obtain an accurate measure of memorization in language models, we would ideally need to compare the model’s output probabilities for every possible sequence against the training data. However, given the vast space of possible sequences, this is computationally intractable. Therefore, researchers have developed various methods to approximate and quantify memorization.
III-A Exposure metric
The Exposure Metric, introduced by Carlini et al. [30], provides a quantitative measure of how much a model has memorized specific sequences from its training data. This metric is particularly useful for assessing the memorization of rare or unique information.
The exposure metric has been widely adopted in memorization studies. To practically apply the exposure metric, researchers often use the “canary extraction test” [30]. This involves inserting known ‘canary’ sequences into the training data and then measuring their exposure in the trained model.
Helali et al. [48] propose the so-called “d-exposure” metric as a measure of memorization for discriminative tasks such as text classification, sinse in those situations we don’t have access to the perplexity of a given text.
III-B Inference attacks
@@
These attacks aim to determine whether a specific data point was part of the model’s training set. Shokri et al. [49] introduced this concept for machine learning models, and it has since been adapted for language models. While the MIA approach in the context of machine learning uses a “shadow learning” paradigm, this is not feasible for LLMs, since they are much heavier in size. Instead, methods for MIA on LLMs try to extract the certainty of the LLMs on the given text and use it as sign of membership in the training data, which is usually done by computing the perplexity metric on the sequence [50, 51, 52, 53, 54, 55, 56].
@@
In conclusion, while measuring memorization in language models remains a challenging task, techniques like the Exposure Metric and Inference Attacks provide valuable insights into the memorization behavior of these models. These methods allow researchers to approximate the extent of memorization and assess potential privacy risks associated with large language models.
III-C Counterfactual memorization
Previous work analyzed the memorization of large language models on sensitive information (e.g. phone numbers) in the training data [13] or synthetically injected ‘canaries’ [30, 60]. However, not all the memorized texts are equally interesting. Zhang et al. [61] dig into a new kind of memorization which is counterfactual memorization. The idea is to see how the presence or absence of a sample of the dataset, affects the performance of the model on the same sample. This definition of memorization has similarities with the definition of differential privacy. Differential privacy as mitigation strategy will be discussed in Section VI-C. In their experiments, the authors create different subsets of a bigger dataset and then they fine-tune the LM on each of these. Then they consider an item $x$ from the datasets and based on the presence or absence of $x$ in the subsets, they divide the subsets into two groups: IN and OUT. Then they test and report the performance on the IN and OUT group of models by averaging. Their experiments on 400 trained models show the counterfactual memorized data, are generally unconventional text such as all-caps, structured formats (i.e. tables or bullet lists), and multilingual texts. In conclusion, counterfactual memorization provides a novel perspective on understanding what language models retain from their training data.
IV Influences and dynamics of memorization
Understanding the factors that influence memorization in large language models is crucial for developing more efficient, secure, and privacy-preserving systems. This section explores various aspects that affect memorization, from model architecture to training processes.
@@ IV-A Model capacity
The first significant factor influencing memorization is model size. Carlini et al. [23] and Tirumala et al. [31] have both demonstrated that larger models are more prone to memorization and do so more rapidly. This trend persists across different architectures and datasets. Carlini et al. [23] find that the relationship between model size and memorization grows consistently on a log-linear scale, with larger models memorizing a greater portion of the data.
Tirumala et al. [31] further highlight that larger models not only memorize more but do so faster in the training process. Interestingly, while memorization increases with model size, it does not necessarily correlate with improved performance. This was shown by Carlini et al. [23] by comparison of models with similar capacities but differing performance levels because of their architectures. Overall, the findings suggest that the ability of LLMs to memorize is strongly linked to their size, potentially due to the high capacity of these models to store detailed information from training data.
@@ IV-B Training data characteristics
The nature of the training data heavily influences memorization. Lee et al. [62] develop tools to deduplicate training data and show that models trained on deduplicated data would produce memorized text ten times less frequently. Kandpal et al. [29] show that a sequence appearing 10 times in the training data is, on average, generated approximately 1000 times more frequently than a sequence that appears only once. A study done by Tirumala et al. [31] on memorization of different parts of speech, reveals that nouns and numbers are memorized significantly faster than other parts of speech, likely because they serve as unique identifiers for specific samples.
@@ IV-C Input and prompting strategies
Carlini et al. [13], McCoy et al. [63], Kandpal et al. [29] show that longer prompts increase the likelihood of triggering memorized sequences, making it easier for language models to regurgitate training data. Moreover, methods like prefix tuning [64] and prompt engineering have been employed to maximize memorization. Ozdayi et al. [65] introduce a novel approach using prompt tuning to control memorized content extraction rates in LLMs. Weller et al. [66] propose ‘according-to’ prompting, a technique that directs LLMs to ground responses in previously observed text.
@@ IV-D Tokenization
Kharitonov et al. [67] explore the impact of the tokenizer on memorization. They experiment with the size of the subword vocabulary learned through Byte-Pair Encoding (BPE) and demonstrate that increasing the subword vocabulary significantly affects the model’s ability and inclination to memorize training data. Furthermore, models with larger vocabularies are more likely to reproduce training data when given specific prompts. The authors suggest that this effect stems from the reduction in sequence lengths as BPE vocabulary size increases.
IV-E Sampling methods
In their experiments, Carlini et al. [13] initially opt for greedy sampling to maximize the regeneration of training data. One limitation is that this sampling scheme generates low diversity outputs; thus, they also experiment with the decaying temperature and Top-n sampling methods, the latter of which shows to be more successful. Yu et al. [68] experiment with different sampling schemes including Decaying temperature, Top-n, Nucleus-$\eta$ and Typical-$\phi$ samplings [69, 70, 71] and use an auto-tuning method on these to find the optimal sampling method that yields to the maximization of training data reproduction.
@@ IV-G Training process dynamics
Zhang et al. [61], Kandpal et al. [29] show that memorization grows consistently with the number of training epochs, which makes sense, as more epochs push the model to potential overfitting. Jagielski et al. [75] demonstrate that examples seen during the earlier stages of training are less prone to memorization and rather they are forgotten over time. These findings indicate that memorization increases with more training, while early-seen examples being more likely to be forgotten.
@@ IV-H Forgetting mechanisms
Kirkpatrick et al. [81] initially introduced catastrophic forgetting in the context of neural networks and continual learning. They propose a method to protect important model weights to retain knowledge. This approach has been effective in maintaining performance on older tasks, even after long periods of non-use. Tirumala et al. [31] observe the forgetting mechanisms of a special batch through the learning process and show that it follows an exponentially degradation, reaching to a constant value baseline. They show that the mentioned baseline scales with the model size. Jagielski et al. [75] address the dual phenomena of memorization and forgetting in LLMs through stronger privacy attacks and several strategies for measuring the worst-case forgetting of the training examples. The study introduces a method to measure to what extent models forget specific training data, highlighting that standard image, speech, and language models do indeed forget examples over time, though non-convex models might retain data indefinitely in the worst case. The findings suggest that examples from early training phases, such as those used in pre-training large models, might enjoy privacy benefits but could disadvantage examples encountered later in training.
Intentional forgetting techniques
As memorization could lead to privacy risks and copyright issues, intentional forgetting could be necessary in some situations. Methods like knowledge unlearning aim to selectively remove specific information from trained models without retraining it from scratch. Bourtoule et al. [82] introduce the SISA framework for efficient machine unlearning, which divides the training data into shards (shards partition the data into disjoint segments) and trains sub-models that can be easily retrained if data needs to be removed. For LLMs specifically, Chen and Yang [83] introduce lightweight unlearning layers into transformers, allowing for selective data removal without full model retraining. Pawelczyk et al. [84] introduce “In-Context Unlearning”, which involves providing specific training instances with flipped labels and additional correctly labeled instances as inputs during inference, effectively removing the targeted information without updating model parameters. Additionally, knowledge unlearning techniques have been categorized into parameter optimization, parameter merging, and in-context learning, each offering unique advantages in efficiently removing harmful or undesirable knowledge from LLMs [85]. These methods not only enhance privacy and security but also ensure that the overall performance of the models remains intact, making them scalable and practical for real-world applications [86].
Factor from Section IV Key findings References
Model capacity Larger models memorize more Carlini et al. [23], Tirumala et al. [31]
Training data Duplicated data amplifies memorization Lee et al. [62], Kandpal et al. [29, 29], Tirumala et al. [31]
@@
Input and prompting Longer prompts and prompt tuning can facilitate recall of the memorized suffix. Carlini et al. [13], McCoy et al. [63], Kandpal et al. [29], Li and Liang [64], Ozdayi et al. [65], Weller et al. [66]
Tokenization Bigger tokenizer vocabulary leads to more memorization Kharitonov et al. [67]
Sampling methods While greedy sampling can pinpoint extremely memorized samples, top-n sampling is the most effective method to retrieve more memorized items. Carlini et al. [13], Yu et al. [68]
@@
Fine-tuning The amount of memorization after fine-tuning significantly varies depending on the task. Mireshghallah et al. [87], Zeng et al. [73]
Training process Earlier phases of training are less prone to memorization Zhang et al. [61], Kandpal et al. [29], Jagielski et al. [75]
Forgetting mechanisms Forgetting follows an exponentially decaying curve Tirumala et al. [31], Jagielski et al. [75]
@@
TABLE II: Influences and dynamics of memorization and their key findings (discussed in Section IV)
In conclusion, memorization in LLMs is influenced by a complex interplay of factors ranging from model architecture to training dynamics. Understanding these influences is crucial for developing more robust, privacy-preserving, and ethically sound language models. An overview of the findings for each of the factors discussed in this subsections is provided in Table II.
V Memorization in specific model architectures
While the target of most of the papers discussed in this survey is auto-regressive large language models, in this section we will overview the memorization in other types of language models, as it would help us gain a deeper understanding of this phenomena.
V-A Masked language models (MLM)
Early studies investigating privacy leakage through the memorization of masked language models (MLMs) yielded limited success. For instance, Vakili and Dalianis [88] attempt to extract patients’ medical conditions from BERT models trained on clinical data, concluding that the risk of privacy leakage in BERT-based models is minimal. Similarly, Lehman et al. [89] conduct attacks on BERT models trained on the MIMIC-III corpus, reaching the same conclusion that stronger attacks may be necessary to trigger memorization in MLMs. Jagannatha et al. [90] employ membership inference attacks on clinical MLMs, demonstrating that these models are far less vulnerable to such attacks compared to auto-regressive (generative) models. Carlini et al. [23] extend their previous work on auto-regressive language models by conducting experiments on the T5 [74] model family, trained on the C4 corpus. They demonstrated that scaling laws still apply, but the degree of memorization in MLMs remains significantly lower compared to similarly sized auto-regressive models.
Mireshghallah et al. [87] introduce “likelihood ratio membership inference attacks” and show that privacy risks in MLMs, while smaller, can still be significant if the attacker uses more advanced methods.
V-B Aligned language models
In studies exploring memorization in aligned language models, experiments typically use either targeted prompting or random prompting approaches. However, these methods face two key challenges that limit their effectiveness.
First, the role tag which is prepended to both input queries and model responses disrupts traditional attack methods, such as the prefix prompting methods. Second, alignment training further complicates extraction attempts. Even if attackers could bypass the role tag, the model’s alignment training, designed to ensure ethical responses, reduces the likelihood of disclosing sensitive or inappropriate information.
To address these challenges, Nasr et al. [34] introduce the “Divergence Attack.” This method asks the model to endlessly repeat a specific word, causing it to break from its alignment training and revert to its original language modeling objectives. In some cases, this led to the verbatim reproduction of training data. Notably, this only occurs when single-token words were repeated, as multi-token prompts led to the model continuing the repetition without reproducing training data.
V-C Retrieval augmented generation (RAG)
Retrieval-augmented language models generate text distributions referencing both the parameters of the underlying language model and the information retrieved from a document index. Huang et al. [91] analyze privacy issues in retrieval-based language models. They use the kNN-LM [92] as their target model through API access. They investigate a scenario in which a model creator has a private, domain-specific data-store that improves model performance on domain-specific tasks, but may also contain sensitive information that should not be revealed. They analyze the access to the text completion and perplexity calculation APIs of the model. Their findings show that kNN-LMs are more prone to disclosing private data from their private data-store compared to parametric models. Zeng et al. [93] conduct targeted and prefix attacks on Retrieval-Augmented Generation (RAG) frameworks to examine how their memorization behaviors differ from those of standard LLMs. The results showed that incorporating retrieval mechanisms significantly lowers the model’s memorization capacity, reducing the success rate of training data extraction attacks in both targeted and prefix attack scenarios. This suggests that the use of retrieval-based data in RAG models makes them more resilient to such attacks compared to standalone LLMs, as the external retrieval component limits the model’s reliance on internal memorization.
VI Mitigating memorization: strategies and techniques

### v2

@@ III Measuring memorization
Measuring memorization in generative language models was originally performed using the exact match metric [47]. However, as we will discuss in Section III-A this has its own limits and issues, therefore we discuss different methods that can be used to provide an approximation of the memorization of a specific model through so-called white-box attacks.
III-A String match
Measuring verbatim memorization (Section Definition) with the methods discussed in the retrievability section (Section II-B), requires exhaustively interacting with the model by inputting different prompts and comparing the model output to the training data. Then the attack success rate is measured by dividing the portion of memorized text by the size of the training data. Since there are infinite combinations of tokens that one can feed the model, this method usually falls short in providing the accurate amount of memorization, rather it can provide a good approximation on memorization lower bound. Since this is directly linked to the memorization risks, it is predominantly used by the researchers [47, 40]. As the exact match metric is too sensitive to small perturbations in the generation, approximate match (Section Definition) has been introduced [26] to capture more memorized samples by bypassing negligible changes in the outputs.
III-B Exposure metric
Introduced by Carlini et al. [24], exposure provides a quantitative measure of how much a model has memorized specific sequences from its training data. This metric is particularly useful for assessing the memorization of rare or unique information.
Exposure is computed based on the negative log-rank of the generation of a sequence according to a language model. To practically apply the exposure metric, researchers often use the “canary extraction test” [24]. This involves inserting known secret numbers as ‘canary’ sequences into the training data and then measuring their exposure in the trained model.
Definition (Exposure).
Given a canary $s[r]$, a model with parameters $\theta$, and the randomness space R, the exposure of $s[r]$ is [47]:
$exposure_{\theta}(s[r])=log_{2}|R|-log_{2}rank_{\theta}(s[r])$ (1)
Helali et al. [58] propose the so-called “d-exposure” metric as a measure of memorization for discriminative tasks such as text classification, since in those situations we don’t have access to the perplexity of a given text.
III-C Inference attacks
@@
These attacks aim to determine whether a specific data point was part of the model’s training set. Shokri et al. [59] introduced this concept for machine learning models, and it has since been adapted for language models. Doing MIA on LLMs typically involves determining the model’s confidence in a given text and using it as an indicator of whether the text was part of the training data. This is often done by computing the perplexity of the sequence [60, 61, 62, 63, 64, 65, 66].
@@
III-D Counterfactuality
Previous work analyzed the memorization of large language models on sensitive information (e.g. phone numbers) in the training data [13] or synthetically injected ‘canaries’ [24, 70]. However, not all the memorized texts are equally interesting. Zhang et al. [71] propose another measure of memorization which is counterfactual memorization. The idea is to see how the presence or absence of a sample of the dataset, affects the performance of the model on the same sample. This measure has similarities with the definition of differential privacy. [^7] Differential privacy as a mitigation strategy will be discussed in Section V-B.
In their experiments, the authors create different subsets of a bigger dataset and then they fine-tune the LM on each of these. Then they consider an item $x$ (e.g. a document from Wikipedia as the dataset) from the datasets and based on the presence or absence of $x$ in the subsets, they divide the subsets into two groups: IN and OUT. Then they test and report the performance on the IN and OUT group of models by averaging. Their experiments on 400 trained models show the counterfactually memorized data, are generally unconventional texts such as all-caps, structured formats (i.e. tables or bullet lists), and multilingual texts. In conclusion, counterfactuality could be used as a metric for measuring memorization, however, the interplay between this metric and the different types of memorization is unexplored and worth more research in the future.
III-E Prompt Compression
Schwarzschild et al. [72] introduce the Adversarial Compression Ratio (ACR) as a novel metric to assess memorization in large language models (LLMs). ACR evaluates whether a string from the training data is memorized by determining if it can be elicited with a significantly shorter adversarial prompt. The method uses the GCG algorithm [73] to find the most compressed prompt. This approach aligns conceptually with Kolmogorov complexity [^8] Kolmogorov complexity measures the shortest possible set of instructions (or program) needed for a computer (Turing machine) to generate a specific output., as it measures the minimum description length of a string but adapts the concept for practical use in LLMs by focusing on adversarial prompts. ACR provides a flexible and compute-efficient framework to evaluate memorization for arbitrary strings, particularly in scenarios like monitoring unlearning and ensuring compliance with data usage policies. This metric offers practical and legal insights into potential misuse of training data, addressing critical concerns in LLM development.
Answer to RQ2. The most dominant method used to measure memorization of LLMs is the string match, however, inference attacks, exposure metrics, counterfactuality, and prompt compression methods can also be used to measure memorization in a more informative way.
IV Influencing factors and dynamics of memorization
Understanding the factors that influence memorization in LLMs is crucial for developing more efficient, secure, and privacy-preserving systems. This section explores various aspects that affect memorization, from model architecture to training processes.
@@ IV-A Model capacity
The first significant factor influencing memorization is model size. Carlini et al. [39], Tirumala et al. [25], Kiyomaru et al. [74], demonstrated that larger models are more prone to memorization and do so more rapidly. This trend persists across different architectures and datasets. Carlini et al. [39] find that the relationship between model size and memorization grows consistently on a log-linear scale, with larger models memorizing a greater portion of the data.
Tirumala et al. [25] further highlight that larger models not only memorize more but do so faster in the training process. Interestingly, while memorization increases with model size, it does not necessarily correlate with improved performance. This was shown by Carlini et al. [39] by comparison of models with similar capacities but differing performance levels because of their architectures.
In the same line, several works analyze the memorization capacity of Transformers from a theoretical perspective. Mahdavi et al. [75] explore the memorization abilities of multi-head attention mechanisms, showing that the number of memorized sequences scales with the number of heads and context size under specific assumptions about input data. Similarly, Kim et al. [76] quantify the theoretical lower bound of memorization capacity of Transformers on sequence-to-sequence mappings and examine this capacity across classification and language modeling tasks with empirical validation.
Overall, the findings suggest that the ability of LLMs to memorize is strongly linked to their size, potentially due to the high capacity of these models to store detailed information from training data.
@@ IV-B Training data characteristics
The nature of the training data heavily influences memorization. Lee et al. [77] develop tools to deduplicate training data and show that models trained on deduplicated data would produce memorized text ten times less frequently. Kandpal et al. [23] show that a sequence appearing 10 times in the training data is, on average, generated approximately 1000 times more frequently than a sequence that appears only once. A study done by Tirumala et al. [25] on memorization of different parts of speech, reveals that nouns and numbers are memorized significantly faster than other parts of speech, likely because they serve as unique identifiers for specific samples. Prashanth et al. [78] categorize memorization into recitation (memorization of highly duplicated sequences), reconstruction (generation using learned templates or patterns), and recollection (memorization of rare, non-template sequences). They argue that memorization is a multi-faceted phenomenon influenced by duplication, template patterns, and rarity, requiring nuanced analysis.
As simple sequences, such as repeated patterns or numbers, are easily memorized by models but often lack substantive content or sensitive information, distinguishing the memorization of these trivial sequences from more complex and meaningful ones would be essential. Duan et al. [28] observed that data with lower z-complexity [^9] Z-complexity is a way to measure how much temporary memory (workspace) a computer needs to generate a specific output using the shortest possible instructions. Unlike Kolmogorov complexity, which only looks at the length of the instructions, Z-complexity also considers the memory used during the process. leads to faster decreases in training loss, as more compressible patterns are memorized more quickly. They further demonstrated that strings of varying complexity exhibit distinct memorization curves, with lower-complexity strings being memorized more easily even for smaller repeats, following a log-linear relationship in memorization probability.
@@ IV-C Input and prompting strategies
Carlini et al. [13], McCoy et al. [79], Kandpal et al. [23] show that longer prompts increase the likelihood of triggering memorized sequences, making it easier for language models to regurgitate training data. Moreover, methods like prefix tuning [80] and prompt engineering have been employed to maximize memorization. Ozdayi et al. [81] introduce a novel approach using prompt tuning to control memorized content extraction rates in LLMs. Wang et al. [82] introduce a dynamic, prefix-dependent soft prompt approach to elicit more memorization. By generating soft prompts based on input variations, this method outperforms previous techniques in extracting memorized content across both text and code generation tasks. Kassem et al. [83] investigate how instruction-tuning influences memorization in LLMs, showing that instruction-tuned models can reveal as much or more training data as base models. They propose a black-box prompt optimization method where an attacker LLM generates instruction-based prompts, achieving higher memorization levels than direct prompting approaches. Weller et al. [84] propose ‘according-to’ prompting, a technique that directs LLMs to ground responses in previously observed text.
@@ IV-D Tokenization
Kharitonov et al. [85] explore the impact of the tokenizer on memorization. They experiment with the size of the sub-word vocabulary learned through Byte-Pair Encoding (BPE) and demonstrate that increasing the sub-word vocabulary significantly affects the model’s ability and inclination to memorize training data. Furthermore, models with larger vocabularies are more likely to reproduce training data when given specific prompts. The authors suggest that this effect stems from the reduction in sequence lengths as BPE vocabulary size increases.
IV-E Decoding methods
While memorization phenomenon stems from the model internals, decoding methods have an important role in arousing it. In their experiments, Carlini et al. [13] initially opt for greedy decoding to maximize the regeneration of training data. One limitation is that this decoding scheme generates low-diversity outputs; thus, they also experiment with the decaying temperature and Top-n decoding methods, the latter of which shows to be more successful. Yu et al. [86] experiment with different decoding schemes, including decaying temperature, top-n, nucleus-$\eta$, and typical-$\phi$ decoding [87, 88, 89] and use an auto-tuning method on these to find the optimal decoding method that yields to the maximization of training data reproduction. As could be drawn from table III, most of the previous works use greedy decoding to heighten memorized content generation, suggesting it could serve as a standard approach for future researchers as well.
@@ IV-G Training process dynamics
Zhang et al. [71], Kandpal et al. [23] show that memorization grows consistently with the number of training epochs, which makes sense, as more epochs push the model to potential overfitting. Jagielski et al. [93] demonstrate that examples seen during the earlier stages of training are less prone to memorization and rather they are forgotten over time. These findings indicate that memorization increases with more training, while early-seen examples being more likely to be forgotten. Leybzon and Kervadec [94] show higher memorization rates happens early and late in training, with lower rates mid-training. Therefore, suggesting that placing sensitive data in the middle stages of training could reduce its vulnerability to extraction attacks.
@@ IV-H Forgetting mechanisms
Kirkpatrick et al. [100] initially introduced “catastrophic forgetting” in the context of neural networks and continual learning. They propose a method to protect important model weights to retain knowledge. This approach has been effective in maintaining performance on older tasks, even after long periods of non-use. Tirumala et al. [25] observe the forgetting mechanisms of a special batch through the learning process and show that it follows an exponential degradation, reaching a constant value baseline. They show that the mentioned baseline scales with the model size. Jagielski et al. [93] address the dual phenomena of memorization and forgetting in LLMs through stronger privacy attacks and several strategies for measuring the worst-case forgetting of the training examples. The study introduces a method to measure to what extent models forget specific training data, highlighting that standard image, speech, and language models do indeed forget examples over time, though non-convex models might retain data indefinitely in the worst case. The findings suggest that examples from early training phases, such as those used in pre-training large models, might enjoy privacy benefits but could disadvantage examples encountered later in training.
IV-I Explainability and interpretability
Huang et al. [101] study verbatim memorization in LLMs using controlled pre-training with injected sequences. They find that memorization requires significant repetition, increases in later checkpoints, and is tied to distributed model states and general language modeling capabilities. Their stress tests show unlearning methods often fail to remove memorized information without degrading model performance, highlighting the difficulty of isolating memorization. Haviv et al. [102] propose a framework to probe how memorized sequences are recalled in transformers, showing that memory recall follows a two-step process: early layers promote the correct token in the output distribution, while upper layers amplify confidence. They find that memorized information is primarily stored and retrieved in early layers. Similarly, Dankers and Titov [103] investigate where memorization occurs across model layers, demonstrating that memorization is a gradual and task-dependent process rather than localized to specific layers. Using centroid analysis, they show that deeper layers contribute more to memorization when models generalize well to new data. Stoehr et al. [104] show that memorization in LLMs, while distributed across layers, is driven by distinct gradients in lower layers and influenced by a low-layer attention head focusing on rare tokens. Perturbation analysis reveals that distinctive tokens in a prefix can corrupt entire continuations, and memorized sequences are harder to unlearn and more robust to corruption than non-memorized ones. Chen et al. [105] reveal interpretability insights into memorization by identifying clustering of sentences with different memorization scores in the embedding space and observing an inverse boundary effect in entropy distributions for memorized and unmemorized sequences. They also demonstrate that hidden states of LLMs can be used to predict unmemorized tokens, shedding light on the dynamics of memorization within the model. These findings offer insights into the mechanisms of memorization, enhancing interpretability and guiding future research.
Factor from Section IV Key findings Representative Works
Model capacity Larger models memorize more [39, 25]
Training data Duplicated data amplifies memorization [77, 23, 25]
@@ IV-I Explainability and interpretability
Input and prompting Longer prompts and prompt tuning can facilitate recall of the memorized suffix. [13, 79, 23, 80, 81, 84]
Tokenization Bigger tokenizer vocabulary leads to more memorization [85]
Decoding methods Greedy decoding is dominantly employed to extract memorized data. [13, 86]
@@
Fine-tuning The amount of memorization after fine-tuning significantly varies depending on the task. [106, 91]
Training process Earlier phases of training are less prone to memorization [71, 23, 93]
Forgetting mechanisms Forgetting follows an exponentially decaying curve [25, 93]
@@
TABLE II: Influencing factors and dynamics of memorization and their key findings (discussed in Section IV)
Work Retrievability Granularity Model Dataset Decoding Prompt Len Match Len
Carlini et al. [39] discoverable verbatim GPT-Neo PILE 50 - 450 50
Tirumala et al. [25] discoverable verbatim roberta wikitext-103 greedy
Borec et al. [107] discoverable verbatim GPT-Neo OpenMemText [107] nucleus 50 50 - 450
Zhou et al. [30] extractable entity GPT-Neo , GPT-J PILE
Wang et al. [82] discoverable verbatim GPT-Neo, GPT-J, Pythia, StarCoderBase PILE, the-stacksmol greedy
Kassem et al. [83] extractable approximate Alpaca, Vicuna, Tulu RedPajama, RefinedWeb, Dolma 66 , 100, 166 133, 200, 366
Leybzon and Kervadec [94] discoverable verbatim OLMo Dolma 32 32
Duan et al. [28] discoverable approximate Pythia-1b, Amber-7b PILE 32 64
Kiyomaru et al. [74] discoverable approximate, verbatim Pythia, LLM-jp PILE, Japanese Wikipedia greedy 100 - 1000 50
Stoehr et al. [104] discoverable verbatim GPT-Neo-125M PILE greedy 50 50
Chen et al. [105] discoverable verbatim Pythia[dedup] PILE greedy 32,48,64,96 32,48,64,96
Huang et al. [101] discoverable verbatim Pythia[dedup] PILE greedy 8, 16, 32, 64 32
Prashanth et al. [78] discoverable verbatim Pythia[dedup] Memorized set of PILE greedy 32 32
Carlini et al. [13] extractable verbatim GPT-2 model generations greedy, temperature 256
Nasr et al. [40] extractable verbatim Pythia, LLaMA, InstructGPT, ChatGPT AuxDataset [40] greedy 50
Shao et al. [108] extractable entity GPT-Neo, GPT-J Enron, LAMA [12] greedy
Huang et al. [109] extractable entity GPT-Neo Enron greedy
Yu et al. [86] discoverable verbatim GPT-Neo-1.3B lm-extraction-benchmark greedy, top-p, top-k, nucleus 50 100
Biderman et al. [42] discoverable verbatim Pythia PILE greedy 32 64
Ippolito et al. [26] discoverable approximate GPT-3, PaLM PILE greedy 50 50
Lee et al. [77] extractable discoverable verbatim T5 [trained again] C4 (variants) top-k >50
Kandpal et al. [23] extractable verbatim Mistral project OpenWebText, C4 top-k, temprature 100 - 700 (char)
Lukas et al. [69] extractable entity GPT-2 ECHR, Enron greedy, beam
Kim et al. [31] extractable entity OPT PILE beam search
Zhang et al. [110] extractable entity GPT-Neo-1.3B PILE greedy, top-p, top-k, beam
Ozdayi et al. [81] discoverable verbatim GPT-Neo lm-extraction-benchmark beam 50 50
TABLE III: Summary and the spectrum of memorization used in different studies: (1) The table highlights that the Pythia and GPT-Neo families are the most commonly used models in memorization studies. This is a logical choice, as they are open-source and available at multiple training checkpoints, making them particularly suitable for analyzing memorization dynamics and contributing factors. Similarly, the PILE dataset is the predominant dataset in these studies, which aligns with the fact that Pythia and GPT-Neo models are primarily trained on it, ensuring consistency in experimental settings. (2) In terms of generation settings, studies largely focus on producing 50-token or 32-token sequences, with greedy decoding being the dominant approach. This preference for greedy decoding is expected, as it maximizes the likelihood of generating memorized sequences, making it an effective choice for probing memorization tendencies. (3) Regarding the spectrum of memorization, approximately 60% of studies focus on discoverability, while the remainder explores extractability. Similarly, verbatim memorization is examined in 60% of works, whereas the rest are split between approximate memorization and entity-level memorization.
Answer to RQ3. In summary, memorization in LLMs is driven by model size, training data properties, input strategies, tokenization, decoding methods, and fine-tuning approaches. Larger models memorize more, duplicated data increases recall, and certain prompts signify memorization. Training dynamics show memorization grows with epochs but can also fade, as models tend to forget earlier samples. Additionally, interpretability research reveals memorization is distributed across layers, with early layers encoding and later layers amplifying content. An overview of the findings for each of the factors discussed in this section is provided in Table II.
V Mitigating memorization: strategies and techniques